What is the formula to figure out the percentage? Example 75=?% of 98..
Answer by spacewolf(2)

When the word "percent" is used, it just means a certain number out of "100" and its u. The general formula for solving percents is the following:


percent = x/100


Here's how you figure out a percent value based on the example you mentioned in your question, "What percent of 98 is 75?":


"98" is your starting number


"75" is the amount out 98 that you have


"x/100" is going to be the percent value


98 * (x/100) = 75 {In math, the general rule is that when you hear the word "of", multiplication is usually involved}


95x = 75*100


x = 7500/95


x = 78.95


Presto: "75" is approximately 79% of "98". Hope this helps.